OpenStudy (anonymous):

Find the intercepts and then graph the line y=4x-2

OpenStudy (anonymous):

x intercept = 1/2 y intercept = -2 so (1/2, -2)

OpenStudy (anonymous):

Can you graph it

OpenStudy (anonymous):

Find the intercepts and then graph the line y=-4x

OpenStudy (anonymous):

x intercept = 0 y intercept = 0

OpenStudy (anonymous):

btw dont write (1/2, -2) as an answer from before, just write it like x intercept = 1/2 or y intercept = -2, or you can do (1/2,0) and (0,-2) but never together, their 2 different poitns their suppose to be